You've observed the following returns on Blast It Corporation's stock over the past five years: 19 percent, -23 percent, 31 percent, 18 percent, and -7 percent, respectively. What was the variance of the returns over this period?

must be entered as decimals rather than percentages, with 4 digits to the right of the decimal point.

**I'm seeing contradicting answers on here, some are saying 438.8% and some are saying .04838 and I don't know which one is correct, if any**
Answer #1

Average return=Total return/Total time period

=(19-23+31+18-7)/5=7.6%

Return (Return-Average return)^2
19 (19-7.6)^2=129.96
-23 (-23-7.6)^2=936.36
31 (31-7.6)^2=547.56
18 (18-7.6)^2=108.16
-7 (-7-7.6)^2=213.16
Total=1935.2%

Standard deviation=[Total (Return-Average return)^2/(Time period-1)]^(1/2)

=[1935.2/(5-1)]^(1/2)

=[1935.2/4]^(1/2)

=22%(Approx)

Variance=Standard deviation^2

=0.0484(Approx).